OpenStudy (anonymous):

Find all the whole number values of n, that would make the following statement true. (3n+9)/(n+1)

OpenStudy (mayankdevnani):

means answer should be 0

OpenStudy (mayankdevnani):

@Mello

OpenStudy (anonymous):

Thanks, but there was a set of numbers on the example question. 0 works, what else?

OpenStudy (mayankdevnani):

(3n+9)/(n+1) it is wholly divided....

OpenStudy (anonymous):

What do you mean by, wholly divided?

OpenStudy (mayankdevnani):

means the remainder is zero(0)

OpenStudy (anonymous):

Right. But I'm pretty sure there are other values of n, that also leaves a remainder of 0

OpenStudy (chihiroasleaf):

n = 2 ?

OpenStudy (mayankdevnani):

no one whole no. can be fully satisfied with n

OpenStudy (mayankdevnani):

@Mello

OpenStudy (mayankdevnani):

there would be some integers..

OpenStudy (chihiroasleaf):

I think \[n = 3k-1, k \in \mathbb{N}\]

OpenStudy (anonymous):

Oh I got it, 0; 1; 2; 5; -2; -3; -4; -7 Heh, thanks for your help :)

OpenStudy (mayankdevnani):

thanx for who??

OpenStudy (mayankdevnani):

@Mello

OpenStudy (mayankdevnani):

are you there??

OpenStudy (chihiroasleaf):

if n is whole number, so negative integers can't be the solution, since whole number starts at 0,1,2,3,....

OpenStudy (anonymous):

\[\frac{3n+9}{n+1}=\frac{3n+3+6}{n+1}=3+\frac{6}{n+1}\]so \(n+1|6\) and we have\[n+1=\pm1,\pm2,\pm3,\pm6\]

OpenStudy (chihiroasleaf):

@mukushla are negative numbers also the solutions? since the question ask for n to be whole number

OpenStudy (anonymous):

u r right negatives are not solution

OpenStudy (anonymous):

only solutions are \(n=0,1,2,5\)

OpenStudy (anonymous):

@mayankdevnani everyone who helped. @chihiroasleaf @mukushla Sorry, I translated this from another language. I think the correct term was integer. Thanks again to everyone for your input!

OpenStudy (anonymous):

np :)